Cereal boxes are filled at a factory. Through extensive sampling, it is found that their equipment fills the boxes to a mean of 502 grams, with a standard deviation of 3.1 grams.

10,000 boxes of cereal are produced each day. 

In a daily run, how many cereal boxes are expected to weigh between 500 and 504 grams?



1
Expert's answer
2021-12-20T16:42:53-0500

Let "X=" the weight of cereal box: "X\sim N(\mu, \sigma^2)"

Given "\mu=502\ g, \sigma=3.1\ g"


"P(500<X<504)=P(X<504)-P(X\leq 500)"

"=P(Z<\dfrac{504-502}{3.1})-P(Z<\dfrac{500-502}{3.1})"

"\approx P(Z<0.64516)-P(Z<-0.64516)"

"\approx0.7405887-0.2594113"

"\approx0.481177"

"10000(0.481177)=4812"

In a daily run, 4812 cereal boxes are expected to weigh between 500 and 504 grams.